# Grounding Fallacies Misrepresenting Scientific Publications in Evidence

### v1

@@ 5 Retrieving Relevant Passages
The passages linked to the reasoning gaps (via $s_{i}$) or to the accurate premise ($p_{0}$) are needed for the fallacious argument reconstruction and are essential for effective debunking, which must explain why a claim was thought to be true and why it actually is not (Lewandowsky et al., 2020).
5.1 Finding the Kernel of Truth
closed (annotated) open (all)
@@ 5.1 Finding the Kernel of Truth
Table 1: Finding a passage $S_{j}^{0}$ based on which the claim was made (a) among the annotated passages only (closed) or (b) across the entire publication (open). Averaged over five seeds (three for LLMs).
Given an incorrect claim $\overline{c}$ that misrepresents a publication $S=[S_{0},S_{1},...,S_{n}]$, the model must rank all passages $S_{i}\in S$ such that the top-ranked passage communicates the information ($p_{0}$) based on which the false claim $\overline{c}$ was made (denoted as $S_{j}^{0}$). This task is similar to finding supporting evidence passages in automated fact-checking (Thorne et al., 2018), but it differs in that the evidence passage is in a “corrupted” support relationship with the claim, meaning it only supports the claim when a fallacy is involved. Finding the passage $S_{j}^{0}$, which explains the basis of the claim, is crucial as it reveals the (broken) rationale behind the claim. We report P@$1$ and MRR over the subset of annotated passages with comprehensive annotations (closed) and over all passages (open). The open evaluation is more realistic, but only serves as a lower bound.
5.1.1 Models
As baselines, we randomly shuffled passages (random) or maintained their original order from the publication (ordered). We used BM25 for lexical similarity-based ranking. For semantic similarity, we ranked the passages by their cosine similarity to the claim using sentence embeddings from BioBERT (Lee et al., 2020), fine-tuned by Deka et al. (2022) for evidence selection in scientific AFC, and prompt-based embeddings from INSTRUCTOR (Su et al., 2022). We also report the performance of the IMS (Wright et al., 2022) used during the dataset construction. For cross-encoders, we trained DeBERTaV3 (He et al., 2022) models on three scientific AFC datasets SciFact (Wadden et al., 2020), CovidFact (Saakyan et al., 2021), HealthVer (Sarrouti et al., 2021), and their union (all) (cf. §D). Given the claim and the passage, we compute and rank passages based on the predicted label probability for Supported. This label fits closest to the task definition of finding a passage that seemingly supports the claim. We used LLMs only in the closed evaluation due to their computational cost. Here, we implemented PRP (Qin et al., 2023), which reorders passages through pairwise comparisons akin to the early iterations of the bubble sort algorithm. We used GPT-4 (OpenAI, 2023) and GPT-3.5333https://platform.openai.com/docs/models/ as commercial LLMs, Llama3-8B [^4] https://llama.meta.com/llama3/ and Llama2-70B (Touvron et al., 2023) as open-source LLMs. The implementation details are outlined in §E.
5.1.2 Results
@@
Table 1 shows solid performance across all evaluated models. The strong ordered baseline suggests that claims often rely on early parts of a study. In the open evaluation, embedding-based approaches perform the best. Note that the IMS model preselected the passages for annotation (in a simplified setup that provided the accurate premise instead of the claim), and its performance must be interpreted with caution. AFC models are superior to embedding models in the closed evaluation, but fall behind PRP ranking via LLMs.
5.2 Finding Undermining Passages
@@
Given an incorrect claim $\overline{c}$ that misrepresents a publication $S=[S_{0},S_{1},...,S_{n}]$, and a passage $S_{j}^{0}\in S$, based on which the claim was made, the model must rank all passages $S_{i}\in S$ such that the top-ranked passages $S_{i}$ are linked to reasoning gaps (expressed via the publication context $s_{i}$) between $S$ and $\overline{c}$. This passage ranking task starkly differs from evidence retrieval in AFC because the model must reason about the relationships between the claim, the accurate premise ($p_{0}$) on which the claim is based, and the content of the passage. For example, in vitro experiments with cell cultures only indicate a fallacy because the claim relies on the results of these experiments ($p_{0}$) but incorrectly suggests that they indicate a cure for humans. Additionally, this task seeks content that reveals a fallacy by disrupting the reasoning, which is different from retrieval for multi-hop reasoning that connects multiple pieces of information to establish a reasoning path. Since multiple passages are often needed to find all fallacies, we report the passage-level mean average precision (MAP) as our main metric. We further report P@$1$ because one detectable fallacy is the minimum requirement to reject a claim. To measure how many distinct reasoning gaps can be detected, we report the fallacy-level recall of the ten top-ranked passages (Fall-R@$10$). This penalizes models that only detect passages linked to the same $s_{i}$. We only evaluate on the open subset, as most annotated passages are linked with reasoning gaps as per the dataset construction.
@@
5.2.1 Models
We use the ranking models from §5.1. We slightly adapted the AFC-based ranking to use the sum of the predicted probabilities for the labels Supported and Refuted. We further assume oracle outputs from the previous subtask and prepended a randomly sampled gold $S_{j}^{0}$ to the claim in all baselines. These design choices follow our experiments on the validation split (cf. §F).
5.2.2 Results
@@
Table 2: Retrieving passages linked to fallacies from the entire publication (open). Averaged over five seeds.
To solve this task, the model must first understand the (false) rationale behind the claim as expressed in the passage $S_{j}^{0}$ (e.g., communicating that “chloroquine reduced infection of the coronavirus”). Then, the model can identify how a different passage interacts with this reasoning to highlight a reasoning gap (e.g., that this was observed in “in vitro” experiments). Table 2 shows that cosine-similarity-based ranking outperforms all AFC-based models that can jointly reason over the claim and the evidence despite the complex reasoning that is required. This suggests that lexical or semantic similarities correlate sufficiently strongly with passages that indicate reasoning gaps, while AFC models seemingly fail to reason over them.
To understand whether fallacy classes impact the difficulty of passage retrieval, we show the recall of detectable fallacy classes over the number of retrieved passages for the INSTRUCTOR and AFC (all) in Figure 3. Overall, within the same model, different fallacy classes follow similar trends. The superior performance of INSTRUCTOR seemingly relies on fallacies that can be detected from $S_{j}^{0}$ passages. For comparison, we visualize the performance without $S_{j}^{0}$ passages in §F.2, Figure 9, which reduces the performance gap between the two ranking models.
6 Fallacious Argument Reconstruction
Given an incorrect claim $\overline{c}$ and relevant selected passages $S_{j}\in S$ of the misrepresented study, the model must verbalize the fallacious premises $\hat{\overline{p}}_{i}$ with the fallacy class $\hat{f}_{i}$ such that all reasoning gaps between $S$ and $\overline{c}$ can be bridged via the generated fallacious premises $\hat{\overline{P}}=[\hat{\overline{p}}_{1},\hat{\overline{p}}_{2},...,\hat{\overline{p}}_{n}]$. We assume oracle output of both passage retrieval tasks and only consider passages linked to a fallacy, ensuring that at least one passage is linked to the accurate premise (as the basis for the claim). We experimented with GPT-4 Turbo as a proprietary LLM, GPT-3.5 as a lower-cost alternative, and Llama3-8B as a representative small open-source LLM with strong leaderboard performance. As in Missci, we prompted the LLMs in a zero-shot setting (cf. §G.1 for prompt selection). Figure 4 shows examples for the Definition, Logical Form and toy Examples for the Fallacy of Composition taken from literature (Bennett, 2012; Cook et al., 2018). By default, all prompts include the these properties as supplementary information for each fallacy (cf. §A for all fallacy classes).
@@ 6 Fallacious Argument Reconstruction
LLM Passages R@5 P@1 Arg@1
Llama3-8B per-passage 0.226 0.290 0.476
Llama3-8B all passages 0.199 0.266 0.425
GPT-3.5 per-passage 0.165 0.190 0.361
GPT-3.5 all passages 0.089 0.194 0.206
6.1 Holistic Argument Evaluation
The evaluation on MissciPlus faces two challenges: First, in Missci, each reasoning gap was expressed by one publication context ($s_{i}$) for which dedicated fallacious premises exist. LLMs were prompted for each $s_{i}$ separately, and the predicted fallacious could be directly compared to the annotated fallacies for the same $s_{i}$. MissciPlus does not have a direct mapping between the evidence passage $S_{j}$ and the dedicated fallacies since the same passage can be linked to multiple reasoning gaps ($s_{i}$) and vice versa. Second, the arguments can contain different reasoning gaps ($s_{i}$) that can be bridged with different fallacious premises that apply the same fallacy class. For example, the claim that “hydroxychloroquine is a cure for COVID-19” in Figure 1 falsely assumes (among others) that i) hydroxychloroquine will have an equal effect as chloroquine, and that ii) SARS-CoV-2 will behave equally to SARS-CoV-1, both of which use the False Equivalence fallacy, but address distinct gaps. This further complicates matching the generated fallacies to the predicted ones. To address this, we evaluated the fallacious argument holistically: Given all relevant passages of the misrepresented study, we expect up to five ranked fallacies ($\hat{\overline{p}}_{i}$, $\hat{f}_{i}$), where five is the maximum number of distinct reasoning gaps in Missci. Following (Schlichtkrull et al., 2023), we automatically match the generated fallacies to the gold fallacies at the argument level (instead of using $s_{i}$ as in Missci).
Evaluation and Matching
Given a predicted fallacy $\hat{y}_{i}$ = ($\hat{\overline{p}}_{i}$, $\hat{f}_{i}$) and a gold fallacy $y_{k}$ = ($\overline{p}_{k}$, $f_{k}$), we define a function $\phi:(y,\hat{y})\rightarrow\{0,1\}$ which discerns if the predicted and the gold fallacy are a match based on two implementations:
The $\phi^{\mathrm{f}}$ is an upper bound, as it does not penalize models for poorly phrased premises. $\phi^{\mathrm{f+p}}$ uses a Llama3-8B model fine-tuned with QLoRA adapters (Dettmers et al., 2023) on the human evaluation data from Missci (cf. §G.2). The performance via cross-validation is 79.8 points in accuracy (78.8 in F1-macro), suggesting that it is suitable, albeit imperfect, in matching the fallacious premises. We do not perform a human evaluation, which is too complex given the many-to-many relationship between the predicted and the gold fallacies, and is hard to reproduce for future work. Instead, we report evaluation measures based on the two complementary implementations $\phi$, which we deem adequate to answer our research questions. As a primary measure, we report the recall of reasoning gaps for which $\phi$ found a match among the five fallacy predictions (R@$5$). Following Glockner et al. (2024a), we use precision P@$1$, which checks whether the top-ranked fallacy is correct based on $\phi$, and Arg@$1$, which considers an argument as successfully rejected if at least one of the predicted fallacies is correct based on $\phi$.
@@ 6.2 Experiments
6.2.1 Benefit of Per-Passage prompting
We compare the performance when prompting LLMs based on each passage $S_{j}$ individually (per-passage) and when including all concatenated passages in a single prompt (all passages). The per-passage prompts follow Missci, where the fallacies where detected for each publication context ($s_{i}$) with the provided accurate premise ($p_{0}$). To apply this on MissciPlus, we replaced the paraphrased publication context $s_{i}$ with a passage $S_{j}$ that is linked to $s_{i}$, and replace $p_{0}$ with a randomly sampled passage $S_{j}^{0}$. For both prompting strategies, we selected the five top-ranked fallacy predictions per argument for evaluation using $\phi^{\mathrm{f+p}}$ in Table 4.
Per-passage prompting is superior for both LLMs, but more expensive due to multiple prompts per argument. The higher (or similar) precision suggests that focusing on one passage (per-passage prompting) can be advantageous when identifying the fallacies. Llama3-8B outperforms GPT-3.5 across all measures, but detects none of the annotated fallacies for more than half of the misinformation, according to the automated metrics.
6.2.2 Missci vs. MissciPlus
Table 3 compares the performance of all three LLMs and ablations over various fallacy information in the prompts (Definition, Logical Form and Explanation) on Missci and MissciPlus using per-passage (per-publication context $s_{i}$ on Missci) prompting. We can see that GPT-4, as a state-of-the-art LLM, performs best across all measures and datasets on this task. We further observe a considerable impact of different fallacy information for each LLM. Yet, none is universally beneficial or harmful for all tested LLMs.
@@ 7 Scientific AFC Evaluation
AFC models rely on retrieved evidence to refute misinformation (Guo et al., 2022). However, evidence that directly refutes a claim may not exist (Glockner et al., 2022), and misinformation often rather distorts the facts (Brennen et al., 2020). MissciPlus contains paired claims with real-world evidence passages that can be assessed by fallacy detection models and evidence-based fact-checking models. Using MissciPlus, we now test how well AFC models and LLMs can detect the nuanced real-world misrepresentation of scientific biomedical publications.

### v2

@@ 5 Retrieving Relevant Passages
A prerequisite to reconstruct the fallacious argument is the identification of the relevant passages in the study. These passages provide evidence why the claim was made, and why this involves fallacies. They are needed for the fallacious argument reconstruction and for effective debunking, which must explain why a claim was thought to be true and why it actually is not (Lewandowsky et al., 2020).
5.1 Subtask 1: Finding the Kernel of Truth
Given an incorrect claim $\overline{c}$ that misrepresents a publication $S=[S_{0},S_{1},...,S_{n}]$, the model must rank all passages $S_{i}\in S$ such that the top-ranked passage communicates the accurate premise ($p_{0}$) based on which the claim $\overline{c}$ was made (denoted as $S_{j}^{0}$). This task is similar to finding supporting evidence in automated fact-checking (Thorne et al., 2018), but differs as the evidence passage is in a “corrupted” support relationship with the claim, meaning it only supports the claim when a fallacy is involved. The passage $S_{j}^{0}$ explains the basis of the claim and reveals the (broken) rationale behind the claim. We report P@$1$ and MRR over the subset of annotated passages with comprehensive annotations (closed) and over all passages (open). The open evaluation is more realistic, but only serves as a lower bound.
As baselines, we randomly shuffled passages (random) or maintained their original order from the publication (ordered). We used BM25 for lexical similarity-based ranking. As semantic embedding based approaches, we ranked the passages by their cosine similarity to the claim using sentence embeddings from BioBERT (Lee et al., 2020), fine-tuned by Deka et al. (2022) for evidence selection in scientific AFC, and prompt-based embeddings from INSTRUCTOR (Su et al., 2023). We also report the performance of the IMS (Wright et al., 2022) used during the dataset construction. Further, we trained DeBERTaV3 (He et al., 2022) AFC models on three scientific AFC datasets SciFact (Wadden et al., 2020), CovidFact (Saakyan et al., 2021), HealthVer (Sarrouti et al., 2021), and their union, denoted as all (cf. §E). Given the claim and the passages, we rank passages based on the predicted label probability for Supported, which fits closest to the task definition of finding a passage that seemingly supports the claim.
Finally, for LLM-based ranking we implemented PRP (Qin et al., 2024), which reorders passages through pairwise comparisons akin to the early iterations of the bubble sort. We used GPT-4 (Achiam et al., 2023) and GPT-3.5 as proprietary LLMs, Llama3-8B (Dubey et al., 2024) and Llama2-70B (Touvron et al., 2023) as open-source LLMs. Implementation details are outlined in §F. Due to the high computational costs of PRP with growing numbers of documents, we only evaluate the LLMs in the closed evaluation with prompt selection based on the development set, Table 17). Preliminary experiments showed cheaper LLM-based methods like list-wise ranking were inefficient due to incorrect outputs and order sensitivity, which is a known limitation (Zhu et al., 2023).
closed open
@@ 5.1 Subtask 1: Finding the Kernel of Truth
Table 1: Finding a passage $S_{j}^{0}$ based on which the claim was made (a) among the annotated passages only (closed) or (b) across the entire publication (open). We list results for (1) baselines, (2) embedding rankers, (3) AFC models and (4) LLMs. Averaged over five seeds (three for LLMs).
@@
Table 1 shows solid performance across all models. The strong ordered baseline suggests that claims often rely on early parts of a study. In the open evaluation, embedding-based approaches perform the best. Note that the IMS model preselected the passages for annotation, and its performance must be interpreted with caution. AFC models are superior to embedding models in the closed evaluation, but fall behind PRP ranking via LLMs.
5.2 Subtask 2: Finding Undermining Passages
@@
Given an incorrect claim $\overline{c}$ that misrepresents a publication $S=[S_{0},S_{1},...,S_{n}]$, and a passage $S_{j}^{0}\in S$, based on which the claim was made, the model must rank all passages $S_{i}\in S$ such that the top-ranked passages $S_{i}$ expose reasoning gaps (i.e., they are linked to the publication context $s_{i}$) between the study content $S$ and the inaccurate claim $\overline{c}$. This passage ranking task differs substantially from evidence retrieval in AFC, as the model must i) understand the rationale behind the claim based on the accurate premise in $S^{0}_{j}$, and ii) evaluate how each passage $S_{i}\in S$ to be ranked impacts this rationale. For example, “in vitro experiments” only indicate a fallacy because the claim relies on the results of these experiments and incorrectly transfers them to humans. This task also differs from multi-hop reasoning (Jiang et al., 2020; Ma et al., 2024), which connects information to establish a reasoning path, e.g., to support or contradict a claim. Instead, the task identifies passages that disrupt the reasoning behind the misinformation.
We report the passage-level mean average precision (MAP) as our main metric. We further report P@$1$ because one detectable fallacy is the minimum requirement to reject a claim. To measure how many distinct reasoning gaps can be detected, we report the fallacy-level recall of the top ten ranked passages (Fall-R@$10$). This penalizes models that only detect passages linked to the same publication context $s_{i}$ and, hence, can only detect the same subset of fallacies. We only evaluate this subtask on the open subset, as most annotated passages are linked with reasoning gaps as per the dataset construction. Here, we slightly adapted the AFC-based ranking to use the sum of the predicted probabilities for the labels Supported and Refuted. We further assume oracle outputs from the previous subtask and prepended a randomly sampled gold $S_{j}^{0}$ to the claim in all baselines. These design choices follow our experiments on the validation split (cf. §G).
To solve this task, the model must first understand the (false) rationale behind the claim as expressed in the passage $S_{j}^{0}$ (e.g., communicating that “chloroquine reduced infection of the coronavirus”). Then, the model can identify how a different passage interacts with this reasoning to highlight a reasoning gap (e.g., that this was observed in “in vitro” experiments). Table 2 shows that cosine-similarity-based ranking outperforms all AFC-based models that can jointly encode the evidence passage with the claim, despite the complexity of the reasoning required. This suggests that lexical or semantic similarities correlate sufficiently strongly with passages that indicate reasoning gaps, while the acquired reasoning by AFC models seem to be not helpful.
@@
Table 2: Retrieving passages linked to fallacies from the entire publication (open) for (1) baselines, (2) embedding rankers and (3) AFC models. Avg over five seeds.
We show the recall of detectable fallacy classes over the number of retrieved passages for the INSTRUCTOR and AFC (all) in Figure 3. Overall, within the same model, different fallacy classes follow similar trends. The superior performance of INSTRUCTOR seemingly relies on fallacies that can be detected from $S_{j}^{0}$ passages, which implies that INSTRUCTOR does not really capture the problematic nature of the passages that undermine the claim. For comparison, we visualize the performance without $S_{j}^{0}$ passages in §G.2, Figure 10, which reduces the performance gap between the two ranking models.
6 Subtask 3: Fallacious Argument Reconstruction
Given an incorrect claim, $\overline{c}$, and relevant selected passages $S_{j}\in S$ from the misrepresented study $S$ that contain the accurate premise $p_{0}$ and the necessary publication contexts $s_{i}$ for detecting fallacies, the model must generate the fallacious premises $\overline{p}_{i}$ along with their corresponding fallacy classes $f_{i}$. These generated fallacious premises, $\hat{\overline{P}}=[\hat{\overline{p}}_{1},\hat{\overline{p}}_{2},...,\hat{\overline{p}}_{n}]$, should bridge all reasoning gaps between $S$ and the incorrect claim $\overline{c}$. We assume oracle output of both passage retrieval tasks from Section 5 and only provide passages linked to a fallacy (with at least one $S_{j}^{0}$ passage as the basis for the claim). We experimented with GPT-4 Turbo and GPT-3.5 as proprietary LLMs , and Llama3-8B as a small open-source LLM with strong leaderboard performance. As in Missci, we prompted the LLMs in a zero-shot setting (cf. §H.1 for prompt selection). By default, all prompts include Definition, Logical Form and toy Examples, as shown in Figure 4, from literature (Bennett, 2012; Cook et al., 2018) as supplementary information for each fallacy (cf. §A for all fallacy classes).
@@ 6 Subtask 3: Fallacious Argument Reconstruction
6.1 Holistic Argument Evaluation
The evaluation on MissciPlus faces two challenges: First, unlike Missci, there is no one-to-one mapping between fallacies and passages. This is because a single passage can simultaneously contain multiple publication contexts ($s_{i}$), each linked to different fallacies, and a single fallacy can also relate to multiple passages. Second, different fallacies ($\overline{p}_{i}$, $f_{i}$) that share the same fallacy class ($f_{i}$) but address different reasoning gaps between the study and the claim may be present. For example, the claim that “hydroxychloroquine is a cure for COVID-19” in Figure 1 is based on two false assumptions: i) that hydroxychloroquine will have the same effect as chloroquine, and ii) that SARS-CoV-2 will behave in the same way as SARS-CoV-1. Each assumption presents a separate error in the claim, both of which are based on the False Equivalence fallacy. Hence, during evaluation, the fallacy class alone is insufficient to determine which of the two problems the model addressed with its generated fallacies. To match the generated fallacies with the gold fallacies, it is essential to additionally use the fallacious premises.
To address these challenges, we evaluated the fallacious argument holistically. Given all relevant passages of the misrepresented study, we expect up to five ranked fallacies ($\hat{\overline{p}}_{i}$, $\hat{f}_{i}$), where five is the maximum number of distinct reasoning gaps, that must be addressed with fallacies, in Missci. Following Schlichtkrull et al. (2023), we automatically match the generated fallacies with the gold fallacies at the argument level (instead of evaluating fallacies per publication context $s_{i}$ as in Missci). We define a function $\phi:(y,\hat{y})\rightarrow\{0,1\}$ which discerns if the predicted fallacy $\hat{y}_{i}$ = ($\hat{\overline{p}}_{i}$, $\hat{f}_{i}$) and the gold fallacy $y_{k}$ = ($\overline{p}_{k}$, $f_{k}$) match based on two implementations:
The $\phi^{\mathrm{f}}$ is an upper bound, as it does not penalize models for poorly phrased premises. $\phi^{\mathrm{f+p}}$ uses a Llama3-8B model, fine-tuned with QLoRA adapters (Dettmers et al., 2023) on the human evaluation data from Missci (cf. §H.2). The accuracy via cross-validation is 79.8% (78.8 in F1-macro). We do not perform a human evaluation, which is too complex given the many-to-many relationship between the predicted and the gold fallacies, and is hard to reproduce for future work. Instead, we report evaluation measures based on the two complementary implementations $\phi$, which we deem adequate to answer our research questions. As a primary measure, we report the recall of reasoning gaps for which $\phi$ found a match among the five fallacy predictions (R@$5$). Following Glockner et al. (2024a), we use precision P@$1$, to check if the top-ranked fallacy is correct, and Arg@$1$, which considers an argument as successfully rejected if at least one of the predicted fallacies is correct.
LLM Passages R@5 P@1 Arg@1
Llama3-8B per-passage 0.226 0.290 0.476
Llama3-8B all passages 0.199 0.266 0.425
GPT-3.5 per-passage 0.165 0.190 0.361
GPT-3.5 all passages 0.089 0.194 0.206
@@ 6.2 Experiments
We compare the performance when prompting LLMs based on each passage $S_{j}$ individually (per-passage) and when including all concatenated passages in a single prompt (all passages). The per-passage prompts follow Missci, but replace the accurate premise ($p_{0}$) and publication context ($s_{i}$) with the respective real-world passages that contain the same information. We always provided oracle passages and selected the top five ranked fallacy predictions per argument for evaluation using $\phi^{\mathrm{f+p}}$ in Table 3. Per-passage prompting shows superior performances for both LLMs, but is more expensive because it requires multiple prompts per argument. The higher (or similar) precision suggests that focusing on one passage (per-passage prompting) can be advantageous when identifying the fallacies. Llama3-8B outperforms GPT-3.5 across all measures, but detects none of the annotated fallacies for more than half of the misinformation, according to Arg@$1$.
Table 4 compares the performance of all three LLMs and ablations over the different fallacy information in the prompts (Definition, Logical Form and Explanation) on Missci and MissciPlus using per-passage prompting (using the paraphrased $s_{i}$ and $p_{0}$ on Missci). GPT-4 performs best across all measures and datasets on this task. We further observe a considerable impact of different fallacy information for each LLM. Yet, none is universally beneficial or harmful for all tested LLMs.
@@ 7 Scientific AFC Evaluation
A key novelty of MissciPlus are the claims paired with real-world evidence passages that can be assessed by fallacy detection models and evidence-based fact-checking models. This allows to test AFC models on fallacious claims.